OpenStudy (anonymous):

Maria has three children. There is two years age difference between each child. The total ages of all three children is 36 years. Tina is the youngest child. How old is Tina? Let T = the age of Tina. Which formula will calculate Tina's age?

OpenStudy (anonymous):

age of tina is t then the age of the elder two of them are t+2 and t+2+2=t+4 respectively as per the problem t+t+2+t+4=36 or 3t+6=36 or 3t=36-6 or 3t=30 or t=30/3 or t=10 thus tina is 10 yrs old